# Thinking Racial Bias in Fair Forgery Detection: Models, Datasets and Evaluations

### v1

@@ 4.1 Bias Offset Problem
where $\mathbb{F}$ donates real face and forgery approaches. $F\in\mathbb{F}$ and $\mathbb{F}=\{\text{RealFace},$
@@
4.2 Aggregation Distortion Problem
Various forgery approaches not only exhibit different fairness situations but also demonstrate distinct levels of performance. For example, in Tab 1, we can see that in Xception, the average accuracy for five forgery methods is 86.24%, 39.47%, 61.36%, 41.67%, and 26.06%, respectively. We identify the aggregation distortion problem where even if we calculate fairness scores separately for each forgery method and average them together, the averaged result can achieve a distorted fairness score due to the performance difference.
Directly averaging fairness scores when employing common fairness metrics might lead to misleading conclusions. For instance, consider two approaches with the accuracy of 20% and 80% respectively. We assume that both approaches yield a bias of 10% if we employ DEO as the fairness metric. Then, we calculate a simple average, which is also 10%. This would lead us to focus solely on the absolute differences in error rates without taking into account the variations in baselines. If the racial biases are both calculated to be 10%, the forgery method with only a 20% accuracy would evidently be much more unfair. This oversimplified average fails to capture the substantial disparity in performance between the two methods. To address the aforementioned issue, we propose a fixed version. For each forgery approach, we have:
@@ 4.2 Aggregation Distortion Problem
where $ACC_{F=f}$ calculates the accuracy of given forgery approach(or real face).
After applying Eq 3 to each forgery method, we calculate their results and then obtain the final fairness score by averaging them. We refer to this approach as Utility Regularized Metric. This nuanced method acknowledges the significance of each forgery approach, providing a more accurate and insightful evaluation of fairness in the context of the diverse fairness and performance landscape. In summary, we recommend not relying on a single fairness metric but rather considering a combination of multiple metrics to collectively reflect the fairness of a detector.
5 Results and Analysis
5.1 Experimental Setup
Algorithms. We summarize the face forgery detection algorithms in Sec A.4. For a comprehensive and fair analysis, we select several representative algorithms. For spatial-based detectors, we choose Xception [1], RECCE [10], UCF [11], Capsule [5], FFD [8] and CORE [9]. For frequency-based detectors, we select F3Net [12], SPSL [13] and SRM [14]. In detail, these models are trained with the Adam optimization algorithm with a learning rate of 0.0002 and an epoch number of 10. The batch size is 32. And data augmentation methods including image compression, horizontal flip and rotation are applied.
Fairness Metric Spatial-based Frequency-based
Xception RECCE UCF Capsule FFD CORE F3Net SPSL SRM
Naive Metric DPD 0.1538 0.1317 0.1782 0.1428 0.1325 0.0914 0.0764 0.0505 0.1335
DEOdds 0.1672 0.1379 0.1655 0.1267 0.1226 0.0858 0.0877 0.0480 0.0982
DEO 0.2095 0.1777 0.2334 0.1849 0.1684 0.1200 0.1030 0.0567 0.1625
STD 0.0563 0.0473 0.0635 0.0513 0.0470 0.0351 0.0278 0.0211 0.0491
Approach Averaged Metric AADPD 0.1957 0.1644 0.2107 0.1655 0.1579 0.1136 0.1102 0.0592 0.1415
AADEOdds 0.1674 0.1378 0.1655 0.1267 0.1255 0.0887 0.0951 0.0512 0.0984
AADEO 0.2099 0.1777 0.2334 0.1849 0.1742 0.1260 0.1178 0.0633 0.1630
AASTD 0.0721 0.0603 0.0773 0.0613 0.0577 0.0442 0.0425 0.0236 0.0523
Utility Regularized Metric URDPD 0.1314 0.1158 0.1433 0.1013 0.1047 0.0805 0.0769 0.0342 0.0917
URDEOdds 0.1069 0.0908 0.1069 0.0782 0.0802 0.0598 0.0624 0.0315 0.0630
URDEO 0.1437 0.1283 0.1614 0.1128 0.1169 0.0909 0.0842 0.0356 0.1060
URSTD 0.0482 0.0423 0.0523 0.0374 0.0381 0.0312 0.0295 0.0136 0.0338
@@ 5.1 Experimental Setup
Table 3: Bias evaluation on FairFD for nine face forgery detectors using Naive Metrics, Approach Averaged Metrics, Utility Regularized Metrics. For each row, we use red to represent the minimum, followed by orange for the second lowest, blue for the third lowest, and dark green for the fourth lowest.
5.2 Benchmarking Fairness of Face Forgery Detectors
The benchmark results (Tab 3) present a comprehensive evaluation of nine face forgery detectors using various fairness metrics. We highlight the four fairest detectors using different colors. Based on the results, we draw the following significant observations:
Current face forgery detectors all exhibit a high degree of racial bias. Even the DPD of SPSL, which has the smallest racial bias, reaches 0.0505. This indicates that the difference in the probability of classifying faces as fake between the most advantaged and disadvantaged groups is 5.05%. When separately calculating and averaging for each forgery method, the AADPD reaches 5.92%. On the other hand, for the least fair detector UCF, the difference in the probability of classifying faces as fake between the most advantaged and disadvantaged groups is 17.82%. This reminds researchers to address the racial bias in face forgery detection.

### v3

@@ Introduction
where $\mathbb{F}$ donates real face and forgery approaches. $f\in\mathbb{F}$ and $\mathbb{F}$ is the set of forgery methods.
Aggregation Distortion Problem
Various forgery approaches not only exhibit different fairness situations but also demonstrate distinct levels of performance. For example, in Figure. 3 and Figure. 4, this is clearly evident. We identify the aggregation distortion problem where even if we calculate fairness scores separately for each forgery method and average them together, the averaged result can achieve a distorted fairness score due to the performance difference.
Directly averaging fairness scores when employing common fairness metrics might lead to misleading conclusions. For instance, consider two approaches with the accuracy of 20% and 80% respectively. We assume that both approaches yield a bias of 10% if we employ DEO as the fairness metric. Then, we calculate a simple average, which is also 10%. This would lead us to focus solely on the absolute differences in error rates without taking into account the variations in baselines. If the racial biases are both calculated to be 10%, the forgery method with only a 20% accuracy would evidently be much more unfair. This oversimplified average fails to capture the substantial disparity in performance between the two methods. To address this issue, we propose a fixed version. For each forgery approach, we have:
URDPD $=\frac{1}{|\mathbb{F}|}\sum_{f\in\mathbb{F}}{\max_{s,s^{\prime}\in\mathbb{S},s\neq s^{\prime}\vphantom{i}}}\bigl{|}P\left(\hat{Y}\,|\,S=s,F=f\right)$
$-P\left(\hat{Y}\,|\,S=s^{\prime},F=f\right)\bigr{|}\enspace/\enspace ACC_{F=f},$ (3)
where $ACC_{F=f}$ calculates the accuracy of given different forgery approaches.
After applying Eq. 3 to each forgery method, we calculate their results and then obtain the final fairness score by averaging them. We refer to this approach as Utility Regularized Metric. This nuanced method acknowledges the significance of each forgery approach, providing a more accurate and insightful evaluation of fairness in the context of the diverse fairness and performance landscape. In summary, we recommend not relying on a single fairness metric but rather considering a combination of multiple metrics to collectively reflect the fairness of a detector.
Bias Pruning with Fair Activations
In this section, we present the Bias Pruning with Fair Activations (BPFA) approach, which develops a novel pruning metric combining weights and the fairness of activations to determine weight importance. Then, we prune those with the lowest pruning scores based on a predefined pruning rate by layer. Here we utilize an unstructured pruning strategy. Noting that the proposed BPFA can be directly extended to process other biases except for racial bias.
Pruning Metric. Consider a convolutional layer weights $W$ of shape $(C_{out},C_{in},S_{Ker}^{h},S_{Ker}^{w})$, where $C_{out}$ represents the number of output filters, and each filter has dimension $(C_{in},S_{Ker}^{h},S_{Ker}^{w})$. For one data sample, the output of this layer is denoted as $X$ with shape $(C_{out},S_{out}^{h},S_{out}^{w})$. The L2 norm by filter of $X$ is represented as $\lVert X\rVert_{2}\in\mathbb{R}^{C_{out}}$. We compute the average L2 norm across all samples from a specific race ${s\in\mathbb{S}}$, denoted by $Z^{s}=\lVert X\rVert_{2}^{s}$. For each filter, we then calculate the standard deviation of these norms across all races, which serves as the bias for that filter:
$BIAS_{i}=std\left(\left\{Z_{i}^{s}\right\}_{s\in\mathbb{S}}\right),$ (4)
where $std(\cdot)$ denotes the standard deviation. This bias measures the variability of outputs across different races. The pruning score (PS) for each weight $W_{ijkm}$ in the convolutional layer at the position ${(i,j,k,m)}$ is then calculated by combining the weight with respect to the computed bias:
$PS_{ijkm}=\frac{\left|W_{ijkm}\right|}{BIAS_{i}}.$ (5)
By comparing the pruning scores, we can identify and potentially remove weights that have the least impact on model utility but the greatest impact on bias. This allows us to reduce bias and improve fairness without sacrificing performance. Note that while our method is illustrated using convolutional layers as an example, it can be easily extended to linear layers.
Benchmark Experiments
Experimental Setup
Algorithms. We summarize the face forgery detection algorithms in the section "Face Forgery Detection Algorithms Categories" in Supp. For a comprehensive and fair analysis, we select several representative algorithms. For spatial-based detectors, we choose Xception [1], RECCE [10], UCF [11], Capsule [5], FFD [8] and CORE [9]. For frequency-based detectors, we select F3Net [12], SPSL [13] and SRM [14]. For fairness-enhanced detectors, we select DAG [37](Xception as base model), DAW [37](Xception as base model) and PFGDFD [38](UCF as base model). In detail, these models are trained with the Adam optimization algorithm with a learning rate of 0.0002 and an epoch number of 10. The batch size is 32. And data augmentation methods including image compression, horizontal flip and rotation are applied. However, when applying these data augmentation methods to DAG, we find that its fairness level significantly deteriorated. For a fair comparison, we report below the results using data augmentation. Meanwhile, the results without data augmentation are presented in the section "Results without Data Augmentation" in Supp.
Fairness Metric Spatial-based Frequency-based Fairness-enhanced
Xception RECCE UCF Capsule FFD CORE F3Net SPSL SRM DAG DAW PFGDFD
@@
[…]
@@
Table 2: Bias evaluation on FairFD for 12 face forgery detectors using Naive Metrics, Approach Averaged Metrics, Utility Regularized Metrics. For each row, the best values are underlined and bolded, followed by the second-best values which are underlined, bolded, and italicized, the third-best values are bolded, and the fourth-best values are bolded and italicized.
Figure 3: Detailed utility (AUC) for diverse races, forgery approaches, detectors.
Benchmarking Fairness of Face Forgery Detectors
[…]
Method DPD DEOdds DEO STD
MA
DPD
MA
DEOdds
MA
DEO
MA
STD
UR
DPD
UR
DEOdds
UR
DEO
UR
STD
AUC ACC
   Original 0.0203 0.0304 0.0215 0.0080 0.0556 0.0481 0.0571 0.0219 0.0320 0.0299 0.0324 0.0126 0.6763 0.7618
   WEIG 0.0183 0.0258 0.0201 0.0072 0.0564 0.0451 0.0586 0.0219 0.0324 0.0277 0.0334 0.0126 0.6769 0.7615
   RoBA 0.1128 0.1598 0.1395 0.0445 0.1462 0.1616 0.1432 0.0583 0.0893 0.1024 0.0867 0.0356 0.6331 0.7037
SPSL    BPFA 0.0181 0.0209 0.0200 0.0072 0.0473 0.0357 0.0496 0.0182 0.0265 0.0218 0.0275 0.0102 0.6862 0.8055
   Original 0.1099 0.1005 0.1242 0.0398 0.1552 0.1196 0.1623 0.0576 0.1037 0.0763 0.1092 0.0384 0.7304 0.5751
   WEIG 0.1098 0.1003 0.1240 0.0398 0.1550 0.1194 0.1621 0.0576 0.1035 0.0761 0.1090 0.0384 0.7304 0.5751
   RoBA - - - - - - - - - - - - 0.5967 -
FFD    BPFA 0.1096 0.0999 0.1237 0.0397 0.1546 0.1189 0.1617 0.0574 0.1032 0.0758 0.1087 0.0382 0.7305 0.5760
   Original 0.0805 0.1396 0.1032 0.0328 0.1393 0.1560 0.1360 0.0530 0.0881 0.0986 0.0860 0.0335 0.6302 0.6019
   WEIG 0.0789 0.1340 0.1012 0.0319 0.1349 0.1494 0.1320 0.0513 0.0853 0.0944 0.0835 0.0324 0.6298 0.6021
   RoBA - - - - - - - - - - - - 0.5468 -
PFG- DFD    BPFA 0.0594 0.1337 0.0796 0.0238 0.1079 0.1442 0.1006 0.0411 0.0644 0.0969 0.0578 0.0245 0.6445 0.7415
Table 3: Experiments with different fairness pruning methods. We highlight the best method for each metric in bold. And we use ’-’ to indicate methods that cause severe performance degradation, rendering the detector unusable even setting a pruning rate as low as 0.1%.